# Exploring a New Design Paradigm for Omnidirectional MAVs for Minimal Actuation and Internal Force Elimination: Theoretical Framework and Control

### v1

@@ IV-B Main Results
where $\zeta_{1e}$, $\dot{\zeta}_{1e}$, $\ddot{\zeta}_{1e}$ $\in se(3)$ are respectively given in (92) and (104). $\bar{K}_{d}$ and $\bar{K}_{a}$ $\in\mathbb{R}^{6\times 6}$ are positive definite diagonal gain matrices chosen such that to make (ABK)(A-BK) Hurwitz where
@@
Given the precompensator in (16) and the linearizing control law in (17), substituting $v$ into (17) asymptotically stabilizes the extended system $\Sigma_{1}$ with (16) in the open neighborhood $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$, where $\mathbb{U}_{\bar{x}_{e}}\subseteq\bar{\mathbb{X}}$ of $\bar{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}\cup\mathbb{D}^{[i]}_{f_{e}}$ is given in (18).
@@
See Sec. VII-D
IV-C Efficient Implementation of the Controller
We introduce here a novel version of the Articulated-body Inertia Algorithm (ABA), originally derived in [24], which is modified to recursively obtain the acceleration $(\ddot{q}_{t},\dot{V}^{0}_{1})$ needed to implement the feedback linearizing law (17) (20) efficiently, instead of inverting the matrix $M$ in (3).
For the system in Fig. 1, a modified ABA computes $(\ddot{q}_{t},\dot{V}^{0}_{1})$ executing the following computations for each $j=2,3,4$
$\begin{split}{M}_{j2}^{A}&={M}^{0}_{B_{j}}\\ \psi_{j2}&=\left(S_{j2}^{T}M^{A}_{j2}S_{j2}\right)^{-1}\\ {W}_{j2}^{A}&=-\operatorname{ad}^{T}_{{V}_{j2}^{0}}{M}^{0}_{B_{j}}{V}^{0}_{j2}-W_{app,B_{j}}^{0}-W_{grav,B_{j}}^{0}\\ {M}_{j1}^{A}&=M_{j2}^{A}-M_{j2}^{A}S_{j2}\psi_{j2}S_{j2}^{T}M_{j2}^{A}\\ \psi_{j1}&=\left(S_{j1}^{T}M^{A}_{j1}S_{j1}\right)^{-1}\\ {W}_{j1}^{A}&=M_{j2}^{A}\left(S_{j2}\ddot{\tilde{q}}_{j2}+\dot{S}_{j2}\dot{q}_{j2}\right)+{W}_{j2}^{A}\end{split}$ (22)
where $\ddot{\tilde{q}}_{ji}$, with $i=1,2$, is given by:
$\begin{split}\ddot{\tilde{q}}_{ji}&=\psi_{ji}\left(\tau_{f_{ji}}-S_{ji}^{T}\left(M^{A}_{ji}\dot{S}_{ji}\dot{q}_{ji}+{W}_{ji}^{A}\right)\right)\end{split}$ (23)
Afterwards, the articulated inertia ${M}_{1}^{A}$ and bias wrench ${W}_{1}^{A}$ for the base are computed as
$\begin{split}{M}_{1}^{A}&={M}^{0}_{B_{1}}+{M}^{0}_{B_{5}}+\sum_{j=2}^{4}\left(M_{j1}^{A}-M_{j1}^{A}S_{j1}\psi_{j1}S_{j1}^{T}M_{j1}^{A}\right),\\ {W}_{1}^{A}&=-\operatorname{ad}^{T}_{{V}_{1}^{0}}({M}^{0}_{B_{1}}+{M}^{0}_{B_{5}}){V}^{0}_{1}-W_{app,B_{1}}^{0}-W_{grav,B_{1}}^{0}-W_{grav,B_{5}}^{0}\\ &\quad+\sum_{j=2}^{4}\big(M_{j1}^{A}(S_{j1}\ddot{\tilde{q}}_{j1}+\dot{S}_{j1}\dot{q}_{j1})+{W}_{j1}^{A}\big).\end{split}$
Hence, the following six linear equations are solved for the base acceleration twist $\dot{V}_{1}^{0}$ using $LDL^{T}$ method, since $M_{1}^{A}$ is symmetric positive definite
$0=M_{1}^{A}\dot{V}_{1}^{0}+{W}_{1}^{A}.$ (25)
The obtained value of $\dot{V}_{1}^{0}$ is then used to retrieve the joint accelerations by evaluating the following expressions:
$\begin{split}\ddot{q}_{j1}&=-\psi_{j1}S_{j1}^{T}M^{A}_{j1}\dot{V}_{1}^{0}+\ddot{\tilde{q}}_{j1}\\ \dot{V}_{j1}^{0}&=\dot{V}_{1}^{0}+S_{j1}\ddot{q}_{j1}+\dot{S}_{j1}\dot{q}_{j1}\\ \ddot{q}_{j2}&=-\psi_{j2}S_{j2}^{T}M^{A}_{j2}\dot{V}_{j1}^{0}+\ddot{\tilde{q}}_{j2}.\end{split}$ (26)
Parameter Value Unit Description
$m_{p}$ $0.5$ kg Link $B_{j}$ mass, $j=2,3,4$
$m_{a}$ $0.5$ kg Link $B_{5}$ mass
$m_{b}$ $1.0$ kg Base $B_{1}$ mass
$a$ $0.5$ m Length between any two adjacent joints
$c$ $0.25$ m Length from a joint to the respective $B_{j}$’s CoM
$g$ $9.81$ m/s2 Gravitational acceleration
$I_{px}$ $2.1875\times 10^{-3}$ kg·m2 Link $B_{j}$ inertia about $x_{p}$-axis
$I_{py}$ $2.1875\times 10^{-3}$ kg·m2 Link $B_{j}$ inertia about $y_{p}$-axis
$I_{pz}$ $1.168\times 10^{-6}$ kg·m2 Link $B_{j}$ inertia about $z_{p}$-axis
$I_{bx}$ $8.3333\times 10^{-3}$ kg·m2 Inertia of base $B_{1}$ about $x_{b}$-axis
$I_{by}$ $8.3333\times 10^{-3}$ kg·m2 Inertia of base $B_{1}$ about $y_{b}$-axis
$I_{bz}$ $1.6667\times 10^{-2}$ kg·m2 Inertia of base $B_{1}$ about $z_{b}$-axis
$b_{fx}$ $0.9$ N.m.s/rad Viscous friction coefficient about $e_{j1}$-axis
$b_{fy}$ $0.9$ N.m.s/rad Viscous friction coefficient about $e_{j2}$-axis
$|k|$ $0.01$ m Ratio of drag-trust coeff. for $B_{j}$’s propeller
TABLE II: Plant Parameters with Definitions and Units
@@ V Simulation Tests
This section presents simulation results designed to establish confidence in the generalizability of our findings beyond the standard model used for formal proofs. To achieve this, we evaluate the proposed controller using a more complex model that incorporates non-idealities such as noise and uncertain parameters. To show that the main body $B_{1}$ of the proposed MAV (4), with the state feedback controller in (17) (20) whose dynamics are given in (16), is capable of stabilizing its attitude independently from its position and vice versa in both designs PJD1 and PJD2, two indicative values for the base pose at equilibrium are selected. The attitude set point is chosen such that the MAV stabilizes a base configuration making the desired frame $\mathcal{F}_{b}$ rotated around the inertial frame $y_{w}$-axis by $60^{\circ}$, i.e., $C^{0}_{1e}=\left(\begin{smallmatrix}R^{0}_{1e}&0_{3\times 1}\\ 0_{1\times 3}&1\end{smallmatrix}\right)$ with $R^{0}_{1e}=R_{x}(0)R_{y}(60^{\circ})R_{z}(0)$, whereas the position step reference is formed by the desired $C_{1e}^{0}$ pose having a constant orientation and a position of $\mathcal{F}_{b}$’s origin equal to 11 mm along the $y_{w}$-axis, i.e., $C^{0}_{1e}=\left(\begin{smallmatrix}I_{3\times 3}&(\,0\,1\,0\,)^{T}\\ 0_{1\times 3}&1\end{smallmatrix}\right)$. The position reference is applied for the first 60 seconds of the simulation. Upon convergence to the desired position, the attitude maneuver starts. The corresponding state and input evolutions over time are reported in Fig. 3.
The orientation and position errors of frame $\mathcal{F}_{b}$ are denoted by $e_{\phi_{1,l}}$ and $e_{p_{1,l}}$, respectively, with $l=\{x,y,z\}$. $e_{\phi_{1,l}}$ is defined as $e_{\phi_{1,l}}:=\phi_{e,l}-\phi_{l}$ where $\phi_{e,l}$ and $\phi_{l}$ are the desired and actual angles resulting from transforming the rotation matrices in $C^{0}_{1e}$ and $C^{0}_{1}$ to the Euler angles sequence $XYZ$ during the simulation time, respectively. $e_{p_{1,l}}$ is given by $e_{p_{1,l}}:=p_{e,l}-p_{l}$ where $p_{e}$ and $p$ are the desired and actual position vectors of $\mathcal{F}_{b}$’s origin in the inertial frame, respectively. The $e_{w^{0}_{1}}$ and $e_{v^{0}_{1}}$ are the evolution of the angular and linear velocities components of the base twist $V^{0}_{1}$, respectively.
To better emulate realistic conditions, the simulation in Fig. 3 includes parametric uncertainty and input noise. The noise is a low-pass filtered Gaussian noise with a state ${\beta}(t)\in\mathbb{R}^{6}$ and dynamics
$\dot{{\beta}}(t)=-k{\beta}(t)+{\mu},$ ${\mu}\in\mathcal{N}({0},l^{2}I_{6\times 6}),$ (27)
where $l=0.1$ and $k=45$. Fig. 6 shows the components of the noise $\beta(t)$ added to the actuators during the simulation in Fig. 3. The true parameter values, listed in Table II, are each perturbed by $5\%$. Additionally, the initial conditions are not an admissible equilibrium and given by $(q_{t}(0),q_{5}(0))=0.1\times\boldsymbol{1}_{8\times 1}\,\text{rad}$, ${x}_{2}(0)=0_{12\times 1}$, and $f(0)=1.05\,f^{[2]}_{e}$, with $f^{[2]}_{e}$ calculated from (49). The initial base configuration for simulation in Fig. 3 is $C^{0}_{1}(0)=\left(\begin{smallmatrix}R^{0}_{1}(0)&(\,1\,-1\,-1\,)^{T}\\ 0_{1\times 3}&1\end{smallmatrix}\right)$ with $R^{0}_{1}(0)=R_{x}(-60^{\circ})R_{y}(-30^{\circ})R_{z}(60^{\circ})$, introducing initial attitude and position errors.
@@
Figure 3: Time evolutions of the state and input for a common simulation are presented for two designs: (a) PJD1 and (b) PJD2. The simulation comprises two phases. First (0-60s), a step reference of 1m along the inertial $y_{w}$ axis is commanded for the base position, with the initial attitude held constant. Second (60-130s, gray-shaded), a step reference demands a $60^{\circ}$ rotation about the inertial $y_{w}$ axis for the attitude, while the position at $t=60$s is maintained. Imperfections such as large initial errors, uncertainty, and noise contribute to the angular velocity and attitude errors during both the initial position maneuver and the subsequent attitude maneuver
Figure 4: Time evolution of the zero dynamics states showing convergence.
Figure 5: Comparison of the total thrust evolution between PJD1 and PJD2 for the same scenario in Fig. 3. In steady state, PJD2 thrust magnitudes sum up to exactly the minimal value (9), unlike its PJD1 counterpart, which is about 2% higher (still much better compared to the internal forces required by state of the art designs).
Figure 6: Evolution of noise states added to the inputs in the simulation in Fig. 3.
To cope with parametric uncertainty, the version of the virtual control including an integral term (108) in Remark 3 is utilized to eliminate the steady state error. The controller gains $\bar{K}_{d}=\operatorname{diag}(5.5576)$ and $\bar{K}_{a}=\operatorname{diag}(14.9433)$ are obtained from applying LQR on the error linear system (107), with weighting matrices on its state $Q=\operatorname{diag}(0.0001)\in\mathbb{R}^{12\times 12}$ and input $R=\operatorname{diag}(0.0001)\in\mathbb{R}^{6\times 6}$, while $K_{I}=\operatorname{diag}(60)$ is selected by trial-and-error to zero the steady-state error. The gain matrix $K_{p}=\operatorname{diag}(0.1)$ appearing in the kinematics controller (92) affects how fast $C^{0}_{1}(t)$ is desired to converge to $C^{0}_{1e}$, which in turn reflects on how fast the reference trajectory $\zeta_{1e}(t)$ in the error system (107) evolves. When the closed-loop error system (107) operates with fixed values of $\bar{K}_{d}$ and $\bar{K}_{a}$, the higher this $K_{p}$ gain is, the larger the error state values become, requiring greater thrusts and faster tilting speed of the actively-tilting propeller.
Furthermore, to show that the zero dynamics (78) in the simulations has a stable behavior as formally proven in Sec. VII-C for the standard model, we simulated its evolution for initial conditions $C^{0}_{1}(0)=I_{4\times 4},\,q_{t}(0)=-30^{\circ}\times\boldsymbol{1}_{6\times 1}\,$, $\dot{q_{t}}(0)=0.1\text{ rad/s }\times\boldsymbol{1}_{6\times 1}$ in the nominal settings. The results in Fig. 4 indicate convergence to zero of the passive joint variables and corresponding speeds, which is indeed the corresponding joint equilibrium in $\bar{x}_{e}$ for the base pose $C_{1e}^{0}=C^{0}_{1}(0)=I_{4\times 4}$.
We conclude by highlighting the main difference between PJD1 and PJD2 by comparing the total thrust magnitudes for the scenario in Fig. 3. It can be observed from Fig. 5 that PJD1 equilibrium thrust satisfies $\sum^{5}_{j=2}f^{[1]}_{je}>\sum^{5}_{j=2}f^{[2]}_{je}$ with $f^{[2]}_{e}=f_{min}\in\mathcal{F}_{\rm min}^{\rm eq}$. In fact, for this desired $C_{1e}^{0}$, the residual torque is $\tau_{dj}\neq 0$ and thus the corresponding joint equilibrium $q_{t}^{[1]}$ yields an equilibrium thrust $f_{e}^{[1]}$ satisfying $\sum^{5}_{j=2}f^{[1]}_{je}\approx 1.02\,f_{min}$, hence witnessing about $2\%$ increase over the minimal value $f_{e}^{[2]}$, which is achieved in PJD2, reinforcing the findings in Proposition 1.
Videos for the numerical simulations in Fig. 3 are available at https://tinyurl.com/bdur26au.

### v2

@@ IV Stabilization of the Equilibria
where $\zeta_{1e}$, $\dot{\zeta}_{1e}$, $\ddot{\zeta}_{1e}$ $\in\mathbb{R}^{6}$ are respectively given in (86) and (98) [29] and shown in Fig. 3. $\bar{K}_{d}$ and $\bar{K}_{a}$ $\in\mathbb{R}^{6\times 6}$ are positive definite diagonal gain matrices chosen such that to make (ABK)(A-BK) Hurwitz where
@@
Given the precompensator in (9) and the linearizing control law in (10), substituting $v$ into (10) asymptotically stabilizes the extended system $\Sigma_{1}$ with (9) in the open neighborhood $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$.
@@
See Sec. -F in [29]. ∎
We note that the system (1) evolves on the product Lie group $SE(3)\times\mathbb{T}^{8}$ due to the internal joint dynamics. This motivates our use of the normal-form approach [27], as it facilitates the study of the associated zero dynamics. Although the proposed controller, shown in Fig. 3, employs tools often considered coordinate-dependent, the Lie-group formulation guarantees that the control laws are coordinate-invariant, i.e. their validity does not depend on a specific local parameterization of the configuration $SE(3)\times\mathbb{T}^{8}$.
@@ V Simulation Tests
This section reports high-fidelity Gazebo simulations, implemented via the GenoM3 framework, validating the proposed controller beyond the nominal model and under more realistic conditions such as presence of parametric uncertainty. For both designs PJD1 and PJD2, the controller (10), (13) with the dynamic extension (9), shown in Fig. 3, achieves independent stabilization of the MAV attitude and position. Two indicative desired base poses are considered: an attitude command of a $60^{\circ}$ step in roll, pitch and yaw at fixed position, and a $1$ [m] position step along the inertial $(x,y,z)$ with constant orientation. The resulting trajectories are shown in Fig. 4. Simulation videos and plots for several different maneuvers are available at https://youtu.be/SWIt0CSwNUo. As noted from these simulations, the vehicle can achieve a wide range of maneuvers while generating only positive thrust, potentially reducing the complexity associated with the use of bidirectional propellers. The state and input evolutions exhibit some oscillations due to the imperfections introduced by large parametric uncertainty and actuator noise from the emulated hardware.
[…]
@@
Figure 4: Evolutions of the state and input for (a) PJD1 and (b) PJD2 designs. The simulation has two phases: First (0-60s) is the position step with the initial attitude held constant. Second (60-120s, gray-shaded), the attitude step with the position at $t=60$s held constant.
Figure 5: zero dynamics evolution showing convergence.
Figure 6: Comparison of the total thrust between PJD1 and PJD2 when stabilizing a $60^{\circ}$ pitch attitude step. In hovering, $\sum f^{[1]}_{e}$ is 2% higher than $\sum f^{[2]}_{e}$.
To illustrate the stability of the zero dynamics (72) [29], as proven in Sec. -E, we simulated its evolution under nominal conditions with initial values $C^{0}_{1}(0)=I_{4\times 4},\,{\color[rgb]{0,0,0}{q_{p}}}(0)=-30^{\circ}\times\boldsymbol{1}_{6\times 1}\,$, $\dot{{\color[rgb]{0,0,0}{q_{p}}}}(0)=0.1\times\boldsymbol{1}_{6\times 1}$ rad/s. As shown in Fig. 5, the $q_{p}$ and $\dot{q}_{p}$ converge to zero, confirming convergence to the equilibrium $q_{pe}$ corresponding to that base pose $C^{0}_{1e}(0)$.
[…]